OpenStudy (anonymous):

Solve for (1/2) (3)² (pi/2)

OpenStudy (unklerhaukus):

\[\frac{1}{2}3^3 \frac{\pi}{ 2} \]

OpenStudy (unklerhaukus):

\[=\frac{9\pi}{4}\]

OpenStudy (anonymous):

i got 2.25

OpenStudy (unklerhaukus):

\[\frac{9\pi}{4}=2.25\pi\]

OpenStudy (anonymous):

ok thank you multily 3.14 with 2.25 and youll get your answer right?

OpenStudy (unklerhaukus):

\[2.25\pi\approx 7.07\]

OpenStudy (anonymous):

thanks. how did you get 7.07 friend?

OpenStudy (unklerhaukus):

just like you said, i multiplied 2.25 by pi or ~3.14 and got 7.06858 i rounded to the hundredths place is 7.07